# Estimators for multivariate allometric regression model

### 17 Feb 2024 v1

@@ 4 Consistency under weak identifiability regimes
    $\bm{X}$ satisfies that $n^{-1}\bm{X}\bm{X}^{\prime}$ converges to a finite positive-definite matrix;
@@
$=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\left(\Theta(n^{2-2\eta})\right)^{2}}=\Theta\left(n^{2\eta-1}\right).$
@@
$\leq\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left(\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right)^{2}}$
@@ 6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are the largest and second largest eigenvalues of $\hat{\bm{\Sigma}}$, respectively. It holds that $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ under some asymptotic regimes such as a traditional large sample regime where $n\to\infty$ with the assumption that $n^{-1}\bm{X}\bm{X}^{\prime}$ converges to a positive-definite matrix while other parameters are fixed. Neither the consistency of $\hat{\bm{\gamma}}_{1}(w^{*})$ nor $\hat{w}^{*}$ necessarily means that a feasible estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is consistent up to sign, that is to say,
@@
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=O(1),$ ( 6.3)
@@
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}+\{{\rm tr}(\bm{\Sigma})-(n-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=o(r_{n}^{2})$
@@
Suppose that, as $n\to\infty$, $n^{-1}\bm{X}\bm{X}^{\prime}$ converges to a positive-definite matrix with other parameters held fixed. Then, as $n\to\infty$, (6.1) holds.

### 24 Feb 2024 v2

@@ 4 Consistency under weak identifiability regimes
    $\bm{X}$ satisfies that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a finite positive-definite matrix;
@@
$\leq\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
@@ 6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are the largest and second largest eigenvalues of $\hat{\bm{\Sigma}}$, respectively. It holds that $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ under some asymptotic regimes such as a traditional large sample regime where $n\to\infty$ with the assumption that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix while other parameters are fixed. Neither the consistency of $\hat{\bm{\gamma}}_{1}(w^{*})$ nor $\hat{w}^{*}$ necessarily means that a feasible estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is consistent up to sign, that is to say,
@@
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=O(1),$ ( 6.3)
@@
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}+\{{\rm tr}(\bm{\Sigma})-(n-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=o(r_{n}^{2})$
@@
Suppose that, as $n\to\infty$, $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix with other parameters held fixed. Then, as $n\to\infty$, (6.1) holds.